OpenStudy (anonymous):

using Pythagorean theorem find b if a=10 and c=16

jimthompson5910 (jim_thompson5910):

a^2+b^2=c^2 a^2+10^2=16^2 a^2+100=256 a^2=256-100 a^2=156 a=sqrt(156) a = 2*sqrt(39)

OpenStudy (anonymous):

i know 10^2=100 and 16^2=256 but 256-100=156 which does not have a perfect square

jimthompson5910 (jim_thompson5910):

oh I swapped a and b, but it's all the same, so b = 2*sqrt(39)

OpenStudy (anonymous):

b^2+100=256, b^2=156, b=sqrt156

OpenStudy (anonymous):

b^2=10^+16^2 =356 b=2sqrt89 (by squaring both sides)

OpenStudy (anonymous):

so the answer is \[2\sqrt{39}\]

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

i am confused on how answer is 2 sqrt 39

OpenStudy (anonymous):

they pulled out a 4 since 4*39=156 4is a perfect square so 2sqrt39

OpenStudy (anonymous):

\[b^2=\sqrt{16^2-10^2}=\sqrt{256-100}=\sqrt{156}=\sqrt{4\times 39}=\sqrt{4}\times \sqrt{39}=2\sqrt{39}\]

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{156}=\sqrt{4*39}=\sqrt{4}\sqrt{39}=2\sqrt{39}\] So \[\large \sqrt{156}=2\sqrt{39}\]

OpenStudy (anonymous):

@jenni, what u want to know? clearify

OpenStudy (anonymous):

what b would equal if a=10^2 and c=16^2 using the Pythagorean theorem

OpenStudy (anonymous):

b would be 2sqrt39 if simplified

OpenStudy (anonymous):

a=10, c=16, b=? a^2+b^2=c^2 a^2=10^2=100 and c^2=16^2=256 b^2=c^2-a^2=256-100=156 b=sqrt[156]=sqrt[4*39]=sqrt[4]*sqrt[39]=2*sqrt[39]